OpenStudy (anonymous):

Which is the area of a circle with a circumference of 15.7 units? (use 3.14 for )

OpenStudy (anonymous):

Here's our equation to find the area of a circle if we know the circumference: A = C^2 / 4π A = 15.7^2 / 4(3.14) A=246.49/12.56 A=19.625